rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#5126 — FS#9051 — Paris Nexus

Attached to Project— ADSL / SDSL
Incident
Backend / Core
CLOSED
100%
The peer-link between th2-11-n5 and gsw-11-n5 is down.
Our customers will go through th2-11-n5.
Date:  Thursday, 01 August 2013, 16:25PM
Reason for closing:  Done
Comment by OVH - Tuesday, 30 July 2013, 02:14AM

We will reboot th2-11-n5,traffic loss expected.


Comment by OVH - Tuesday, 30 July 2013, 06:31AM

We have the same problem between th2-12-n5 and gsw-12-n5.


Comment by OVH - Tuesday, 30 July 2013, 06:31AM

We will reboot gsw-12-n5.


Comment by OVH - Tuesday, 30 July 2013, 06:33AM

gsw-12-n5 is not up following the reboot, we will send a technician to gsw tomorrow morning,currently, our customers will keep going through th2-12.


Comment by OVH - Tuesday, 30 July 2013, 06:34AM

gsw-12-n5 is up now, we will keep monitoring.


Comment by OVH - Tuesday, 30 July 2013, 08:25AM

Now we have a problem on pair-10 gsw-10 and th2-10 , we are reloading th2-10.


Comment by OVH - Tuesday, 30 July 2013, 08:25AM

Everything is OK following the reboot.


Comment by OVH - Tuesday, 30 July 2013, 18:35PM

Due to the optical problem, we lost securisation between
th2-12-n5 and gsw-12-n5.
The customers continue to function on th2-12-n5.


Comment by OVH - Wednesday, 31 July 2013, 01:57AM

We are now facing repetition issues on several optics DWDM equipping our DSL switches in Paris and used to connect our Globalswitch POPs and TH2. This has a several times impact on the traffic (many cutoffs from few seconds to several minutes depending on customers).

We have therefore decided to change the techno and move the links GSW <> TH2 in DWDM active ADVA on a first half the concerned links. This has to allow the stability of the infrastructure and avoid new cutoffs. A team will move tomorrow morning to Paris in order to prepare a new link. The switching will be done the following night on 31/07 to 01/08 from midnight.


Comment by OVH - Thursday, 01 August 2013, 01:19AM

We will start now, we will cut the secondary lines to prepare the switch over.


Comment by OVH - Thursday, 01 August 2013, 06:05AM

We mounted the 1st peer-link on nexus pair-12.


Comment by OVH - Thursday, 01 August 2013, 06:06AM

peer-links are OK,now we are making the uplinks


Comment by OVH - Thursday, 01 August 2013, 06:07AM

peer-link and up-link were all switched on active equipement.


Comment by OVH - Thursday, 01 August 2013, 06:08AM

The new link ADVA is working.